# Towards Passive Safe Reinforcement Learning: A Comparative Study on Contact-rich Robotic Manipulation

### v1

@@ IV-D Real-world experiments
We deploy the trained passivity-aware models in the real world to complete our validation. In real-world experiments as shown in Fig. 8, we implement our method on a Franka Panda robot arm with a flange mounted at the end effector with an ATI45 FT sensor to percept external force and torque, with our passivity-ensured impedance controller (work frequency 1000Hz). We deploy the best models of Eb4, Eb8 and Eb8-Ef7 in the real world without fine-tuning. The robot can finish the task without violating force constraints and energy constraints in each experiment, which validates our method’s effectiveness. However, failure case happened in the passivity-agnostic model as shown in Fig. 8.d. The video is available as supplementary material.
V Conclusion and future work
@@
In this letter, we investigated the safety and stability of the robotic system under multiple constraints for a contact-rich task, demonstrating that the traditional RL approach can not guarantee passivity. Furthermore, we combined an augmented energy tank-based passivity formalism in a safe RL framework. We employed the passivity constraints in both training and deployment of the RL model. The proposed method completed the task without any force and energy constraint violation. Additionally, passivity-aware policies showed improved energy efficiency in deployment, even though they converged later in the training. A set of experiments both in simulation and in the real world on a 7-DoF Franka robot arm validated the effectiveness of the proposed method. We are planning to explore self-adaptive tank-size and flow constraints in the future works.

### v2

@@ III-D Real-world experiments
The trained models are deplyed in the real world to complete our validation on a Franka Panda robot arm with a flange mounted at the end effector with an ATI45 FT sensor to percept external force and torque as shown in Fig. 8 upper, with our passivity-ensured impedance controller (frequency 1000Hz). We run the best models of each setup 3 times in the real world without fine-tuning. The robot can finish the task without violating force constraints and energy constraints in each experiment, which validates our method’s effectiveness. However, failure case happened in the passivity-agnostic model. Specifically, the Fig. 8 lower shows all models’ energy expenditure over time steps. First, we can observe that the agnostic consumes energy most aggressively although keep passive due to passivity layer, while passivity-aware models maintain lower energy consumption rates. Notably, Eb4 achieves the most efficient energy usage among all models. This suggests that training with tighter energy constraints leads to more energy-efficient behaviors that transfer well to real-world deployment. The overall trend validates that our passivity-aware approach successfully learns energy-efficient policies that retain effectiveness in real-world settings. The video is available as supplementary material.
III-E Discussion
During training, the agent becomes more energy-conservative under passivity constraints $C_{e}$. These constraints also reduce task success rates by limiting exploration, especially early in learning, suggesting constrained cases need more training than unconstrained ones. However, overly restrictive constraints can prevent learning entirely if the task becomes physically infeasible.
Deployment results in Table III show that passivity cannot be guaranteed without a passivity layer, regardless of whether the agent was trained with passivity awareness. While the passivity layer improves task completion and reduces violations for all agents, those trained with energy constraints ($C_{e}$) demonstrate better energy efficiency.
The correlation between the energy usage trends in simulation (Fig. 7.a) and real-world experiments (Fig. 8 .d) demonstrates the successful transfer of learned energy-efficient behaviors. In both cases, the agnostic model shows the steepest decline in energy budget, while Eb4 maintains the most conservative energy usage. This consistent pattern validates that the energy-aware policies learned in simulation maintain their characteristics when deployed on physical hardware. The real-world results also confirm that tighter energy constraints during training (Eb4) lead to more efficient policies that better manage the energy budget in both simulated and real environments. However, the absolute rate of energy consumption differs between simulation and reality due to modeling discrepancies and real-world friction effects.
IV Conclusion and future work
@@
In this letter, we investigated the safety and stability of the robotic system under multiple constraints for a contact-rich task, demonstrating that the traditional RL approach can not guarantee passivity. Furthermore, we combined an augmented energy tank-based passivity formalism in a safe RL framework. We employed the passivity constraints in both training and deployment of the RL model. The proposed method completed the task without any force violation while ensuring system passivity.Additionally, passivity-aware policies showed improved energy efficiency in deployment, even though they converged later in the training. A set of experiments both in simulation and in the real world on a 7-DoF Franka robot arm validated the effectiveness of the proposed method. We are planning to explore self-adaptive energy tank-size and flow constraints in the future works.